a flower bed is in the shape of a triangle with one side twice the length of the shortest side, and the third side is 12 feet more than the length of the shortest side. Find the dimensions if the perimeter is 112 feet
Answer by Cromlix(4381)

Hi there,
Make the shortest side = 'x'
One side = 2x
Third side = x + 12
Perimeter is the total of the three sides.
x + 2x + (x + 12) = 112
Collect like terms
4x + 12 = 112
4x = 112 - 12
4x = 100
x = 25 feet.
Shortest side = 25 feet
One side = 50 feet
Third side = 37 feet.
25 + 50 + 37 = 112 feet
Hope this helps :-)